horses are to race over a certain english meadows for a distance of 4.0 furlongs. what is the race distance in [a] rods and [b] chains? [ 1 furlong = 201.168 m. 1 rod = 5.0292 m and 1 chain = 20.117 m]
1
Expert's answer
2012-06-19T11:10:52-0400
4.0 furlongs = 4·201.168 m& = 804.672 m.

[a] in rods

4.0 furlongs = 804.672 m = 804.672/5.0292 rods = 160 rods;

[b] in chains

4.0 furlongs = 804.672 m = 804.672/20.117 chains ≈ 40 chains.